QUESTION 6 If in a given year an Economy produces 200 units of textiles and 300 units of Soybeans, and the price of textiles is $1 and the price of Soybeans is $2, Nominal GDP is equal to while Real GDP is equal to $800; 800 Textiles units $800; 500 textiles units $800; $800 Not enough information
Answer #1

Answer : The answer is option C.

Nominal GDP = Current year price * Quantity = (1 * 200) + (2 * 300) = 200 + 600

=> Nominal GDP = $800

Here Base year and Current year is same. So,

Real GDP = Base year price level * Quantity = (1 * 200) + (2 * 300) = 200 + 600

=> Real GDP = $800

Therefore, option C is correct.